Sanfillipo, Inc., had 800 units of inventory on hand at March 1 of the current year, costing $20 each. Purchases and sales of inventory during the month of March were as follows:

Date Purchases Sales
March 8 600 units
15 400 units @ $22 each
22 400 units @ $24 each
27 400 units


Sanfillipo uses the perpetual inventory system. According to a physical count, 600 units were on hand at the end of March.

The cost of goods sold for March applying the average cost method is:
Answer #1

The average cost is calculated after each transaction under perpetual inventory method

Cost of Goods sold will be as follows:

March 8 Sales = 600*20 = $12,000

March 27 Sales = [(200*20+400*22+400*24)/1000]*400 = $8960

Hence, cost of goods sold for March = $20,960

The answer is $20,960